# Fixed-Time Input-to-State Stability for Singularly Perturbed Systems via Composite Lyapunov Functions

### v1

@@ 1 Introduction
Many complex dynamical systems can be decomposed into components operating on different time scales. Such multi-time-scale systems are common across a diverse array of engineering applications, including aerospace [1], chemical processing [2], power systems [3], smart grids [4], and biological and evolutionary systems [5], among others. Such systems can be studied using techniques from singular perturbation theory [6, 7, 8, 9], where the methods typically focus on decomposing the system into lower order subsystems and studying the subsystems to draw conclusions on the behavior of the overall interconnection.
One simple yet powerful technique for analyzing singularly perturbed systems is the composite Lyapunov method, first introduced by Khalil in [10]. This method is particularly effective for assessing system stability, as it leverages the stability of the subsystems to construct a Lyapunov function for the overall system that is valid for sufficiently small $\varepsilon$. As demonstrated in [10], if the reduced and boundary layer systems admit certain quadratic-type Lyapunov functions, then under additional interconnection conditions, the interconnected system can be shown to be asymptotically stable, provided there is sufficient time-scale separation between the two subsystems. Moreover, if the Lyapunov functions satisfy specific quadratic bounds, these results can be extended to exponential stability. The composite Lyapunov method has found profound applications in control synthesis across various engineering domains, including power systems [11], biological systems [12], aerospace [13], etc.
Fixed-time stability, initially discovered in [14] and later popularized through the introduction of Lyapunov conditions in [15], has garnered significant attention due to its ability to address challenges in control [16], optimization [17], and learning [18, 19]. Fixed-time stability ensures convergence to an equilibrium within a fixed time, regardless of the system’s initial conditions. While the Lyapunov conditions simplify the verification of fixed-time stability in continuous-time dynamical systems, techniques for analyzing fixed-time stability in interconnected systems remain limited. Current state-of-the-art methods are primarily applicable to systems with specific homogeneity properties [20], restrictive structural requirements [21], or, for the slightly weaker notion of finite-time stability, those satisfying a small-gain condition [22].
In many practical applications, the systems of interest are also influenced by external disturbances or exogenous inputs. In particular, in this paper we study systems of the form
@@
In this paper, we address the above questions by extending the composite Lyapunov method [10] to study FxT ISS in nonlinear singularly perturbed systems with inputs. Specifically, under the assumptions that the respective “reduced” and “boundary-layer” dynamics of (1) are fixed-time stable on their own, we derive interconnection conditions under which (1) renders the origin FxT ISS. These interconnection conditions parallel those developed in [10] for asymptotic and exponential stability. However, unlike the results of [10], absence of Lipschitz continuity at the origin results in more complex interconnection conditions that involve powers of the fixed-time Lyapunov functions of the slow and fast dynamics. Additionally, we leverage our analytical tools to synthesize and analyze novel algorithms able to achieve fixed-time feedback optimization under strongly convex cost functions and slowly varying inputs, [28, 29, 30], which, to the best of our knowledge, represents the first result that achieves fixed-time stability properties. Finally, we present numerical and analytical examples to illustrate the key assumptions and the main stability results of the paper.
Additional Contributions with respect to [31]
@@ 2.1 Notation
We use $\mathbb{R}_{\geq 0}$ to denote the set of nonnegative real numbers. We let $\text{sgn}:\mathbb{R}\to\{-1,1\}$ denote the sign function, i.e $\text{sgn}(x)=1$ if $x>0$, $\text{sgn}(x)=-1$ if $x<0$ and $\text{sgn}(0)=0$. Given a continuous function $\alpha:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$, we say $\alpha\in\mathcal{K}$ (i.e, $\alpha$ is of class $\mathcal{K}$) if $\alpha(0)=0$ and $\alpha$ is strictly increasing. If $\alpha\in\mathcal{K}$ also satisfies $\lim_{s\to\infty}\alpha(s)=\infty$, we say $\alpha\in\mathcal{K}_{\infty}$. Given a continuous function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$, we say $\beta\in\mathcal{KL}$ if for each $t\geq 0,$ $\beta(\cdot,t)\in\mathcal{K}$ and for each $r\geq 0$, $\beta(r,\cdot)$ is non-increasing and asymptotically goes to 0. Furthermore, $\beta\in\mathcal{GKL}$ if $\beta(\cdot,0)\in\mathcal{K}$ and for each fixed $r\geq 0$, $\beta(r,\cdot)$ is continuous, non-increasing and there exists some $T_{\beta}(r)\in[0,\infty)$ such that $\beta(r,t)=0$ for all $t\geq T_{\beta}(r)$. The mapping $T_{\beta}$ is called the settling time function. Given a measurable function $u:\mathbb{R}_{\geq 0}\to\mathbb{R}^{n}$ we denote $|u|_{\infty}=\text{ess}\sup_{t\geq 0}|u(t)|$. We use $\mathcal{L}_{\infty}^{n}$ to denote the set of measurable functions $u:\mathbb{R}_{\geq 0}\to\mathbb{R}^{n}$ satisfying $|u|_{\infty}<\infty$. Next, we will present some lemmas that will be instrumental for our results. The following lemma is a direct consequence of Young’s inequality.
@@
Given $x,y\geq 0$, $\underline{\alpha}\in\mathcal{K}_{\infty}$ and $p_{1},p_{2}>0$ there exists $\overline{\alpha}\in\mathcal{K}_{\infty}$ such that the following holds for all $c>0$:
$|x|^{p_{1}}|y|^{p_{2}}\leq\dfrac{1}{\underline{\alpha}(c)}|x|^{p_{1}+p_{2}}+\overline{\alpha}(c)|y|^{p_{1}+p_{2}}.$
@@
Given $x\geq 0$ and $a\leq b\leq c$, then $x^{b}\leq x^{a}+x^{c}$.
$\square$
Lemma 3
For $\xi_{1}\in(0,1)$ and $\xi_{2}<0$, the following hold for all $x,y\in\mathbb{R}^{n}$:
$\left\lvert\dfrac{x}{|x|^{\xi_{1}}}-\dfrac{y+x}{|y+x|^{\xi_{1}}}\right\rvert$ $\leq 2^{\xi_{1}}|y|^{1-\xi_{1}}$
$\left\lvert\dfrac{x}{|x|^{\xi_{2}}}-\dfrac{y+x}{|y+x|^{\xi_{2}}}\right\rvert$ $\leq K_{\xi_{2}}|y|\left(|x|^{-\xi_{2}}+|y|^{-\xi_{2}}\right).$
where $K_{\xi_{2}}:=1+\max\left(1,-\xi_{2}2^{-\xi_{2}-1}\right)$. $\square$
@@ 2.2 Fixed-Time Input-to-State Stability
where $x\in\mathbb{R}^{n}$ is the state and $u\in\mathcal{L}_{\infty}^{m}$ is an input signal. We assume the vector field $f$ is continuous and satisfies $f(0,0)=0$. We will state some results from [25] that will be particularly relevant for our work.
@@
System (3) is said to be fixed-time input-to-state stable (FxT ISS) if for each $x_{0}\in\mathbb{R}^{n}$ and $u(t)\in\mathcal{L}_{\infty}^{m}$, every solution $x(t)$ of (3) exists for $t\geq 0$ and satisfies
$|x(t)|\leq\beta(|x_{0}|,t)+\vartheta(|u|_{\infty}),$
where $\beta\in\mathcal{GKL}$, $\vartheta\in\mathcal{K}$ and the settling time function $T_{\beta}$ of $\beta$ is a continuous and uniformly bounded function with $T_{\beta}(0)=0$.
@@
$\dfrac{\partial V}{\partial x}f(x,u)\leq-k_{1}V^{a_{1}}-k_{2}V^{a_{2}}+\vartheta(|u|),$ (5)
for some $\vartheta\in\mathcal{K}_{\infty}$, $k_{1},k_{2}>0$, $a_{1}\in(0,1)$ and $a_{2}>1$.
It was shown in [25] that if (3) admits a FxT ISS Lyapunov function, then (3) is FxT ISS.
@@
It can be verified using the techniques in [24, Remark 2.4] that the “dissipation” formulation we use in (5) is equivalent to implication (10) in [25]. $\square$
3 Problem Statement
We consider a class of interconnections of the form (1) with states $x\in\mathbb{R}^{n_{1}}$ and $z\in\mathbb{R}^{n_{2}}$, input $u\in\mathbb{R}^{m}$, a small parameter $\varepsilon>0$, and dynamics satisfying $f(0,0)=g(0,0)=0$. We make the following assumption on (1b):
@@ 3 Problem Statement
There exists a $\mathcal{C}^{1}$ mapping $h:\mathbb{R}^{n_{1}}\to\mathbb{R}^{n_{2}}$ such that $g(x,z,0)=0$ if and only if $z=h(x)$. $\square$
This assumption is quite standard in the singular perturbation literature, and the map $h$ is usually referred to as the quasi-steady state [9] for the disturbance-free system. With this, we define the reduced system from (1a):
@@
We can then define the “error” state $y=z-h(x)$ and consider the following system:
@@
System (7b) is studied in the time scale $\tau=t/\varepsilon$, and letting $\varepsilon\to 0^{+}$ to obtain the boundary layer dynamics:
$\dfrac{dy}{d\tau}=g(x,y+h(x),u).$ (8)
We make the following Lyapunov-based stability assumptions on the lower order systems (6) and (8):
@@
For (6), there exists a $\mathcal{C}^{1}$ function $V:\mathbb{R}^{n_{1}}\to\mathbb{R}_{\geq 0}$ and $\alpha_{1},\alpha_{2},\rho_{R}\in\mathcal{K}_{\infty}$ such that
@@
$\dfrac{\partial V}{\partial x}f(x,h(x),u)\leq-k_{1}V^{a_{1}}-k_{2}V^{a_{2}}+\rho_{R}(|u|),$
@@
For (8), there exists a $\mathcal{C}^{1}$ function $W:\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}\to\mathbb{R}_{\geq 0}$ and $\beta_{1},\beta_{2},\rho_{B}\in\mathcal{K}_{\infty}$ such that
$\beta_{1}(|y|)\leq W(x,y)\leq\beta_{2}(|y|),$
@@
$\dfrac{\partial W}{\partial y}g(x,y+h(x),u)\leq-\kappa_{1}W^{b_{1}}-\kappa_{2}W^{b_{2}}+\rho_{B}(|u|),$
@@
Assumptions 2 and 3 are the “fixed-time” counterpart of the assumptions considered in [10]. However, as in [10], the mere stability (ISS) of the reduced and boundary layer systems is usually insufficient to guarantee that (1) is also stable (ISS). As mentioned in [26], in general, it is not possible to draw stability conclusions for the overall system without considering additional interconnection conditions. While for globally Lipschitz systems such conditions are usually easy to verify, for fixed-time stable systems the dynamics of interest are usually not Lipschitz continuous, which makes the analysis more challenging.
3.1 Main Results
To assess the stability of (1) using a Lyapunov-based approach, we need to consider the following interconnection terms:
$I_{1}$ $=\dfrac{\partial V}{\partial x}\left(f(x,y+h(x),u)-f(x,h(x),u)\right)$ (9a)
$I_{2}$ $=\left(\dfrac{\partial W}{\partial x}-\dfrac{\partial W}{\partial y}\dfrac{\partial h}{\partial x}\right)f(x,y+h(x),u).$ (9b)
Ideally, we would like to derive appropriate bounds on these terms that would allow us to conclude that (1) is fixed time ISS. Before we do so, we define the following terms:

### v2

@@ 1 Introduction
Many complex dynamical systems can be decomposed into components operating on different time scales. Such multi-time-scale systems are common across a diverse array of engineering applications, including aerospace [1], chemical processing [2], smart grids [3], and biological and evolutionary systems [4], among others. Such systems can be studied using techniques from singular perturbation theory [5, 6, 7], where the methods typically focus on decomposing the system into lower order subsystems and studying the subsystems to draw conclusions on the behavior of the overall interconnection.
One simple yet powerful technique for analyzing singularly perturbed systems is the composite Lyapunov method, first introduced by Khalil in [8]. This method is particularly effective for assessing system stability, as it leverages the stability of the subsystems to construct a Lyapunov function for the overall system that is valid under sufficiently large time-scale separation. As demonstrated in [8], if the reduced and boundary layer systems admit certain quadratic-type Lyapunov functions, then under additional interconnection conditions, the interconnected system can be shown to be asymptotically stable, provided there is sufficient time-scale separation between the two subsystems. Moreover, if the Lyapunov functions satisfy specific quadratic bounds, these results can be extended to exponential stability. The composite Lyapunov method has found profound applications in control synthesis across various engineering domains, including power systems [9], biological systems [10], aerospace [11], etc. While the composite Lyapunov method has been widely used to establish asymptotic and exponential stability, it has found limited applications in settings which require stronger notions of stability. In [12], the authors leverage composite Lyapunov functions to establish finite-time stability for a special class of singularly perturbed homogeneous systems. However, to the best of the authors’ knowledge, there does not exist a framework that studies fixed-time stability for a general class of singularly perturbed systems via composite Lyapunov techniques.
Fixed-time stability, which has been popularized through the introduction of Lyapunov conditions in [13], has garnered significant attention due to its ability to address challenges in control [14], optimization [15], and learning [16, 17]. Fixed-time stability ensures convergence to an equilibrium within a fixed time, regardless of the system’s initial conditions. While the Lyapunov conditions simplify the verification of fixed-time stability in continuous-time dynamical systems, techniques for analyzing fixed-time stability in interconnected systems remain limited. Current state-of-the-art methods are primarily applicable to systems with specific homogeneity properties [18], restrictive structural requirements [19], or those satisfying certain small-gain conditions [20]. Moreover, in many practical applications, the systems of interest are also influenced by external disturbances or exogenous inputs. In particular, in this paper we study systems of the form
@@
In this paper, we address the above questions by extending the popular composite Lyapunov method [8] to study FxT ISS in singularly perturbed systems with inputs. Specifically, under the assumptions that the respective “reduced” and “boundary-layer” dynamics of (1) are fixed-time stable on their own, we first derive interconnection conditions under which (1) renders the origin FxT ISS. These interconnection conditions parallel those developed in [8] for asymptotic and exponential stability. However, unlike the results of [8], absence of Lipschitz continuity at the origin results in a more complex paradigm for verifying our proposed interconnection conditions, which involve powers of the fixed-time Lyapunov functions of the reduced and boundary-layer dynamics. Subsequently, we employ our analytical tools to design and analyze a novel class of algorithms that achieve fixed-time feedback optimization under strongly convex cost functions and slowly varying inputs—a class of problems widely studied in the literature [26, 27, 28], but which, to the best of our knowledge, has not previously been addressed using algorithms with fixed-time stability properties. Finally, we present numerical and analytical examples to illustrate the key assumptions and the main stability results of the paper.
@@ 2.1 Notation
We use $\mathbb{R}_{\geq 0}$ to denote the set of nonnegative real numbers. We let $\text{sgn}:\mathbb{R}\to\{-1,{0,}1\}$ denote the sign function, i.e $\text{sgn}(x)=1$ if $x>0$, $\text{sgn}(x)=-1$ if $x<0$ and $\text{sgn}(0)=0$. For a continuous function $\alpha:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$, we say $\alpha\in\mathcal{K}$ (i.e, $\alpha$ is of class $\mathcal{K}$) if $\alpha(0)=0$ and $\alpha$ is strictly increasing. If $\alpha\in\mathcal{K}$ also satisfies $\lim_{s\to\infty}\alpha(s)=\infty$, we say $\alpha\in\mathcal{K}_{\infty}$. Given a continuous function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$, we say $\beta\in\mathcal{KL}$ if for each $t\geq 0,$ $\beta(\cdot,t)\in\mathcal{K}$ and for each $r\geq 0$, $\beta(r,\cdot)$ is non-increasing and asymptotically goes to 0. Furthermore, $\beta\in\mathcal{GKL}$ if $\beta(\cdot,0)\in\mathcal{K}$ and for each fixed $r\geq 0$, $\beta(r,\cdot)$ is continuous, non-increasing and there exists a function $T:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ such that $\beta(r,t)=0$ for all $t\geq T(r)$. The mapping $T$ is called a settling time function, which, in general, is not unique. Given a measurable function $u:\mathbb{R}_{\geq 0}\to\mathbb{R}^{p}$ we denote $|u|_{\infty}=\text{ess}\sup_{t\geq 0}|u(t)|$. We use $\mathcal{L}_{\infty}^{p}$ to denote the set of measurable functions $u:\mathbb{R}_{\geq 0}\to\mathbb{R}^{p}$ satisfying $|u|_{\infty}<\infty$. Given a differentiable function $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$, we use $\mathbf{J}_{f}(x)\in\mathbb{R}^{m\times n}$ to denote the Jacobian of $f$ evaluated at $x\in\mathbb{R}^{n}$. If $m=1$, we use $\nabla f(x)=\mathbf{J}_{f}(x)^{\top}$. If $\mathbf{J}_{f}(x)$ is continuous, we say $f$ is $\mathcal{C}^{1}$.
2.2 Auxiliary Results
We first present some lemmas that will be instrumental for our results. Due to space limitations, the proofs can be found in the extended manuscript [30].
@@ 2.2 Auxiliary Results
Given $x,y\geq 0$ and $p_{1},p_{2}>0$, the following inequality holds for all $c>0$:
$|x|^{p_{1}}|y|^{p_{2}}\leq{c}|x|^{p_{1}+p_{2}}+{c^{-\frac{p_{1}}{p_{2}}}}|y|^{p_{1}+p_{2}}.$
@@
For $\xi_{1}\in(0,1)$ and $\xi_{2}<0$, the following inequalities hold for all $x,y\in\mathbb{R}^{n}$:
$\left\lvert\dfrac{x}{|x|^{\xi_{1}}}-\dfrac{y}{|y|^{\xi_{1}}}\right\rvert$ $\leq 2^{\xi_{1}}|x-y|^{1-\xi_{1}}$ (2a)
$\left\lvert\dfrac{x}{|x|^{\xi_{2}}}-\dfrac{y}{|y|^{\xi_{2}}}\right\rvert$ $\leq K|y-x|\left(|x|^{-\xi_{2}}+|y-x|^{-\xi_{2}}\right).$ (2b)
where $K:=1+\max\left(1,-\xi_{2}2^{-\xi_{2}-1}\right)$. $\square$
@@ 2.3 Fixed-Time Input-to-State Stability
where $x\in\mathbb{R}^{n}$ is the state and $u\in\mathcal{L}_{\infty}^{p}$ is an input signal. We assume the vector field $f$ is continuous and satisfies $f(0,0)=0$. We will state some definitions from [23] that will be particularly relevant for our work.
@@
System (3) is said to be fixed-time input-to-state stable (FxT ISS) if for each $x_{0}\in\mathbb{R}^{n}$ and $u\in\mathcal{L}_{\infty}^{p}$, every solution $x(t)$ of (3) exists for $t\geq 0$ and satisfies
$|x(t)|\leq\beta(|x_{0}|,t)+\varrho(|u|_{\infty}),$ (4)
where $\beta\in\mathcal{GKL}$, $\varrho\in\mathcal{K}$ and there exists a settling time function $T$ of $\beta$ that is continuous and uniformly bounded, with $T(0)=0$.
@@
$\dfrac{\partial V}{\partial x}f(x,u)\leq-k_{1}V^{a_{1}}{(x)}-k_{2}V^{a_{2}}{(x)}+\rho(|u|),$ (6)
for some $\rho\in\mathcal{K}_{\infty}$, $k_{1},k_{2}>0$, $a_{1}\in(0,1)$ and $a_{2}>1$.
@@
It can be verified that the “dissipation” formulation we use in (6) implies relation (10) in [23], which implies FxT ISS for (3) via [23, Thm 4]. Indeed, if (6) holds and we fix some $0<\tilde{\varepsilon}<\min_{i}k_{i}$, we have
$|x|>\chi(|u|)\Rightarrow\dfrac{\partial V}{\partial x}f(x,u)\leq-\tilde{k}_{1}V^{a_{1}}{(x)}-\tilde{k}_{2}V^{a_{2}}{(x)},$
where $\chi(\cdot)=\tilde{\chi}^{-1}\circ\rho(\cdot)$, $\tilde{\chi}(\cdot)=\tilde{\varepsilon}\alpha_{2}^{a_{1}}(\cdot)+\tilde{\varepsilon}\alpha_{2}^{a_{2}}(\cdot)$, and $\tilde{k}_{i}=k_{i}-\tilde{\varepsilon}$. By taking $\tilde{\varepsilon}\to 0$, it can be observed via [23, Corollary 2] that systems that admit a FxT ISS Lyapunov function that satisfy (6) also admit a settling time function satisfying the following bound
$T(x_{0})\leq\frac{1}{k_{1}(1-a_{1})}+\frac{1}{k_{2}(a_{2}-1)},$ (7)
for all $x_{0}\in\mathbb{R}^{n}$.
3 Main Results
We consider singularly perturbed systems of the form (1), with states $x\in\mathbb{R}^{n}$ and $z\in\mathbb{R}^{m}$, input $u\in\mathbb{R}^{p}$, dynamics satisfying $f(0,z^{*},0)=g(0,z^{*},0)=0$ for some $z^{*}\in\mathbb{R}^{m}$, and a small parameter $\varepsilon>0$ that induces a time scale separation between the dynamics of $x$ and $z$. Our main objective is to exploit the stability properties of the so-called reduced system and boundary-layer system associated with (1), as defined below, in order to derive Lyapunov-based sufficient conditions that ensure fixed-time input-to-state stability (FxT ISS) of (1), provided the time-scale separation is sufficiently large.
3.1 Assumptions
To study system (1) in the context of singular perturbations, we make the following standard assumption on (1b):
@@ 3.1 Assumptions
There exists a $\mathcal{C}^{1}$ mapping $h:\mathbb{R}^{n}\to\mathbb{R}^{m}$ such that $g(x,z,0)=0$ if and only if $z=h(x)$. $\square$
The map $h$ is usually referred to as the quasi-steady state mapping [7] for the disturbance-free system (1b). By using this mapping, we can define the so-called reduced system from (1a):
@@
Using the change of coordinates y=zh(x)y=z-h(x), we obtain the following error dynamics:
@@
System (9b) is studied in the time scale $\tau=t/\varepsilon$ and taking $\varepsilon\to 0^{+}$ to obtain the boundary layer system:
$\dfrac{dy}{d\tau}=g(x,y+h(x),u),$ (10)
where $x\in\mathbb{R}^{n}$ is considered fixed. Since our goal is to study FxT ISS of (1), we make the following FxT ISS Lyapunov-based assumptions on the lower order systems (8) and (10):
@@
There exists a $\mathcal{C}^{1}$ function $V:\mathbb{R}^{n}\to\mathbb{R}_{\geq 0}$ and $\alpha_{1},\alpha_{2},\rho_{R}\in\mathcal{K}_{\infty}$ such that
@@
$\dfrac{\partial V}{\partial x}f(x,h(x),u)\leq-k_{1}V^{a_{1}}{(x)}-k_{2}V^{a_{2}}{(x)}+\rho_{R}(|u|),$
@@
There exists a $\mathcal{C}^{1}$ function $W:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}_{\geq 0}$ and $\tilde{\alpha}_{1},\tilde{\alpha}_{2},\rho_{B}\in\mathcal{K}_{\infty}$ such that
$\tilde{\alpha}_{1}(|y|)\leq W(x,y)\leq\tilde{\alpha}_{2}(|y|),$
@@
$\dfrac{\partial W}{\partial y}g(x,y+h(x),u)\leq-\kappa_{1}W^{b_{1}}{(x,y)}-\kappa_{2}W^{b_{2}}{(x,y)}+\rho_{B}(|u|),$
@@
Assumptions 2 and 3 state, respectively, that the reduced system (8) and the boundary-layer system (10) admit individual FxT ISS Lyapunov functions and are, hence, FxT ISS. These assumptions mirror the classical Lyapunov-based conditions commonly used in the literature to study the asymptotic stability of singularly perturbed systems. However, as shown in [8] and [24], the mere stability or ISS properties of the individual reduced and boundary-layer systems are usually insufficient to guarantee that (1) is also stable or ISS, and additional interconnection conditions need to be examined.
3.2 Analysis
To assess the FxT stability of system (1) using a Lyapunov-based approach, we consider the following Lyapunov function candidate
$\Psi_{\zeta}(x,y)=\zeta V(x)+(1-\zeta)W(x,y),\quad\zeta\in(0,1),$ (11)
where $V$ and $W$ are from Assumptions 2 and 3 respectively. For historical reasons, we refer to (11) as a composite Lyapunov function [31, 32]. Evaluating the Lie derivative of (11) along the trajectories of (9) results in
$\dot{\Psi}_{\zeta}$ $=\zeta\left(\frac{\partial V}{\partial x}f(x,h(x),u)+I_{1}(x,y,u)\right)$
$~~~+(1-\zeta)\left(\frac{1}{\varepsilon}\frac{\partial W}{\partial y}g(x,y+h(x),u)+I_{2}(x,y,u)\right),$ (12)
where the interconnection terms, $I_{1}$ and $I_{2}$, are given by
$I_{1}{(x,y,u)}$ $=\dfrac{\partial V}{\partial x}\left(f(x,y+h(x),u)-f(x,h(x),u)\right)$ (13a)
$I_{2}{(x,y,u)}$ $=\left(\dfrac{\partial W}{\partial x}-\dfrac{\partial W}{\partial y}\dfrac{\partial h}{\partial x}\right)f(x,y+h(x),u).$ (13b)
Ideally, we aim to derive suitable bounds on these terms that would allow us to conclude that (1) is fixed-time ISS. Before we do so, we define the following terms: